OpenStudy (anonymous):

The perimeter of a rectangle is 30 ft. The length is 5 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting systems. Let x be the length and y be the width. Write first equation. 2x + 2y= _______ Write Second equation. x=y+_____ What is the length?______ft What is the width?_______ft

OpenStudy (zzr0ck3r):

2x+2y = 30 x = y+5 where x is length

OpenStudy (zzr0ck3r):

2x+2y = 30 x-y = 5 now multiply the second one by -2 system is now 2x+2y = 30 -2x+2y = -10 add them together 4y = 20 y = 5 x = y+5 = 5+5 = 10

OpenStudy (zzr0ck3r):

make sense @WrittenOnMyHeart ?

OpenStudy (anonymous):

Yeah, it does. The length is 10 and the width is 5.

OpenStudy (zzr0ck3r):

yes but do you see how I got there?

OpenStudy (anonymous):

Yes. I see it's the very first step that always gets me. Guessing a working number or whatever you call it.